LUCI/LBC Technical Observing, prep for Partner Science:
UT 2010 Dec 28-30

Observer: 28 Dec: RGreen (@LBT) DThompson, DLMiller, ARakich (Tucson)
Telescope Operator: DGHuerta
Instrument Support:
Telescope Support:
SW support:

Summary

Mostly bad weather forecast for these nights, so combining the logs...

Preparations

Details

28 Dec Closed dome testing of LUCI

LUCI1 functional checkout:
  • Confirmed dates/times are the same on lucifer and obs3
  • showSupervisorStatus lists all OK
  • MCE/ROE powered, confirmed on Web-IO Instrument Control
  • Restarted TelescopeService because IIF was restarted (yesterday?)
    everything else looks OK.
  • UI started, looks OK
  • Moved mask from No Mask to #3 Closed/blind. Took 3:06 minutes
  • Darks, 2sec O2DCR (files 2,3) 11.50e-, 10s MER (files 4,5) 5.40e- readnoise
  • Moved mask from closed/blind to Spectro Sieve. Took 6:06 minutes
  • Sieve mask images to check DIQ:
    • N3.75 (0008) 2.9 pixels FWHM
    • N1.8 (0010, dark=0011) center @ 1.8 pixels FWHM, top (3.4x2.3 pix ovals) & bottom (1.7 pix w/tails) obviously aberrated
  • Move calibration unit in
    • All lamps work
    • All gratings complete tilt setup
  • Check for red leaks in NB filters...
    • 0024: N1.8 200_H+K/1.93 clear+OrderSep Halo2
    • 0025 Br_gam - 0037 HeI (2 images on H2, 10 & 100s)
    • All look good, no obvious thermal leaks
  • LUCI1 flexure tests: /home/LBTO/Flexures/RotatorScript.txt
    • N3.75, mirror, Spectro Sieve, K clear filters, 2s o2dcr
    • luci_20101228_0002 - 0031 (0001 to check levels)
    • mirror_N375_20101228.ps sent to Walter and attached to IT#3046

*Old data deleted from /data/luci through end of Nov. Now 22GB free space on disk.

29 Dec LBC Science Prep

TCS BP7 was running, switching to AO5 for LBC
Could not get firefox to come up under LBTO account (critical IT#)
LBC: Blue channel CCD Controller error:
  • 01:46:54.929000 CAMERA CCDCTRL "application buffer full" (code:21) controller error [ccdctrl.c:1528]
  • 01:46:54.925013 CAMERA CCD Controller error, errcode:8 [src/camera/camera.c:556]
  • Olga says to "Turn OFF" and then "Turn ON"...seems to have worked.

10Bias_Bino ran OK
02:22 WT10_310 on old pointing model...switching to LBCDX.20101215.ptmod in PCSInstrument.conf
PCS restarted to load new pointing model
+1000nm Z11 on SX added after switch to TCS AO5.
FPIA trying to converge but taking forever.
3:15 or so, stopped FPIA because we hit a force limit on DX M1.
  • Revert to start of night (clear active optics, c00, instr offsets)
  • Cannot clear an error in RX on the DX mirror, calling Chris

04:15 Finally working on RED. DX within 1 degree, SX about 2 degrees warmer than outside, both dropping. Still having trouble with large negative Z22 being driven in by DOFPIA causing bad fits in blue even though we start with a decent looking pupil.

05:05 Changed "blue Z22 zero" from 1.6 to 2.3, restarted IDL, re-run dofpia. Seems to converge.

Note added later by Olga Kuhn: See wiki log from 2010 April 08 UT at http://wiki.lbto.org/bin/view/PartnerObserving/20100408logUT for details on the change of BlueZ22Zero from 2.3 to 1.6

05:15 Run science-type script to check guiding (littledumbell), works OK.

05:31 dofpia again here to refocus, then run superfoc

06:37 (EL~81) lbcrangebal 063442 063437 send+ test-
  • didn't find collimation files
  • copied to Sandbox
  • loaded Observe package for moveOPE
  • we hit a limit on DX we should not have
  • not clear what happened and out of time for LBC.

07:02 Reconfigure for LUCI + BP7

29 Dec testing of LUCI

Created new rb-LUCI collimation models:
  • SXPMLUCIFERCollimation.dat and SXSMLUCIFERCollimation.dat
  • Also made dated copied tagged 20101229.
  • Restarted PSF but then realized problem was the X connection was frozen

Correct pointing (0003)

Pointing model:

/home/telescope/PointingLogs
  • 20101229_SX.log
  • 20101229_DX.log
ID RA---J2000---DEC LUCI# ID RA---J2000---DEC LUCI#
ACT0372 09 10 13.230 +54 58 43.13 0010 WT10_401 15 37 45.834 +71 09 13.25 0011
ACT0395 03 49 38.224 +67 09 55.79 0012 WT10_357 05 48 14.529 +46 43 18.28 0013
WT10_280 02 54 17.555 +20 33 56.95 0014 ACT0195 05 15 23.693 -03 21 57.48 0015
ACT0269 07 14 50.691 +18 33 42.91 0016 WT10_095 06 46 13.627 -34 14 34.20 0017
WT10_184 09 48 12.704 -07 14 44.66 0018 WT10_305 09 36 48.963 +27 45 30.64 0021
WT10_224 12 29 28.162 +04 50 17.65 0019 ACT0324 12 21 11.138 +34 41 12.05 0020
WT10_337 11 19 41.501 +38 06 3.15 0022 WT10_292 11 55 24.510 +25 31 20.01 0023
ACT0258 11 33 56.775 +13 33 7.28 0024 WT10_288 08 55 27.406 +26 24 35.35 0025
WT10_201 09 08 1.539 -02 28 34.25 0026 WT10_069 08 48 54.566 -38 58 12.35 0027
ACT0129 08 01 44.532 -21 09 34.63 0028 WT10_303 08 00 32.129 +29 12 44.47 0029
ACT0250 06 07 54.597 +12 38 49.05 0030 WT10_229 04 03 21.222 +03 11 16.18 0031
ACT0282 04 26 18.500 +21 28 13.55 0032 ACT0333 07 51 6.087 +37 12 42.26 0033
ACT0397 07 25 10.536 +67 08 49.81 0034 ACT0405 23 55 15.647 +74 52 25.83 0035
WT10_409 11 10 55.137 +81 43 54.00 0036 ACT0348 09 04 27.247 +43 26 30.93 0037
ACT0273 10 09 8.435 +18 31 50.12 0038      
  • LUCI center of rotation 1013.14 1042.73

0054 reasonable collimation on BS star at 86EL
0055 ACT0224 62EL
0056 WT10_135 36EL
0057 ACT0188 EL53
0058 ACT0224 EL63
0059? 0060 WT10_277 EL75
0061 ACT0309 EL86

Observing Plan (copy and paste this to the nightly logs)

LUCI1 functional checkout:
  • Check system time is same as our GPS ntp server
  • Run /home/engineer/bin/showSupervisorStatus.sh
  • Check LUCI1 Web-IO Control, verify that ROE and MCE are powered
  • Start LUCIFER Control SW and check that all is running as needed
  • Open UI and check that telescope data are being updated
  • Check darks (see: LUCIFER$/readnoise script)
    • O2DCR 2.0s (usually around 11.57 e-)
    • MER10 10.0s
  • Take sieve mask images in N3.75 and N1.8 to check focus gradients
  • Move calibration unit in
    • Check that lamps work
    • Check gratings/tilts
    • Take an imaging flat and check levels
    • Move calibration unit out
  • LUCI1 flexure tests (Nancy's script)

Other Tests:

  • Closed dome SW tests, if requested

  • Twilight flats in NB, BB filters (cycle through all over time)
    • Need to set up a records page on the wiki to keep track

  • Pupil imaging, several rotations (two EL?)

  • Correct initial collimation/pointing

  • Open dome SW tests, if requested
    • Optical pointing/co-pointing/range balancing algorithm, also using IRTC

  • Test any new template observing scripts prior to release

  • Run science-type dice5 pattern to check that things still work
    • Preset to available BS91XX star from UI/Catalog
    • Execute script: ~/LBTO/SciencePrep/dice5_10x (~1 hr)
    • Calculate rms/position, scale & rotation of pattern

  • Repeat check of dithering along slit (20x 2' dither)
    • Preset to available BS91XX star from UI/Catalog
    • Execute script: ~/LBTO/SciencePrep/slitdither.acq
    • Align BS source to center of 2" longslit
    • Execute script: ~/LBTO/SciencePrep/slitdither.sci (~30 min)
    • Calculate AGw alignment to slit

  • Test focus task and verify it does what is needed on LUCI.

  • Repeat measurement of focal plane scale, orientation

  • Additional work on Andrew's M1/M2 collimation?

  • Check existing collimation model
    • Take new data if necessary
    • Stable conditions!

  • Check existing pointing model
    • Implement new collimation model first, if necessary
    • Stable conditions!

  • Take a set of (monocular) range-balanced collimation data
    • Stable conditions!
    • implement this collimation model

  • Take a set of pointing data using range-balanced collimation model
    • Stable conditions!

  • Collect field aberration data

  • Photometric zero point and extinction monitoring

  • Collect a "movie" of 2s LUCIFER images while an in-field offset is taking place.
    • Check to make sure offsets still happen in a reasonable manner

  • Collect through-focus sequence for Andrew - reasonably dense star field.
    • Try Z4 of -1000, -500. 0, 500, 1000 nm
    • Also check the AGw-LUCIFER focus in cold weather for temp term?

  • Collect astrometric field data to check AGw calibration for monitoring

  • Improve map of AGw patrol field limits at high field angle
    • Requires set up work on a stone field

  • Gather NB illumination correction data (BrG done on 20100606)

  • Re-test updatePointingReference functionality.

  • Continue gaining experience with LUCIFER on-sky (e.g. spec acq.)

  • Test dithering between extreme slits on MOS mask + telluric
    • Manual needs 0.59836mm/" scale for the dXdY calculation
    • The how-to set up telluric observations needs updating for clarity.

  • Investigate variable imaging FF gradients by taking data while the telescope is offset with respect to the enclosure (will need manual overriding of normal operations) to see if the observed gradients could be caused by something 'external to the instrument'.

-- DavidThompson - 28 Dec 2010
Topic revision: r10 - 30 Oct 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback